Three fair dices are rolled together. First dice Shows odd number, second dice shows 1 or 3 and whether third dice shows a 5 or 6- check the events are. independent or not.


1
Expert's answer
2021-06-15T07:53:11-0400

There are "6^3=216" outcomes for rolling three fair dices together.

Suitable outcomes


"(1, 1, 5), (1, 1, 6), (1, 3, 5), (1, 3, 6),"

"(3, 1, 5), (3, 1, 6), (3, 3, 5), (3, 3, 6),"

"(5, 1, 5), (5, 1, 6), (5, 3, 5), (5, 3, 6)"

The probability of the event "odd number on first and 1 or 3 on second, and 5 or 6 on third" is


"P(I\cap II\cap III)=\dfrac{12}{216}=\dfrac{1}{18}"

The probability of the event "odd number on first" is


"P(I)=\dfrac{3}{6}=\dfrac{1}{2}"

The probability of the event "1 or 3 on second" is


"P(II)=\dfrac{2}{6}=\dfrac{1}{3}"

The probability of the event "5 or 6 on third" is


"P(III)=\dfrac{2}{6}=\dfrac{1}{3}"

Hence


"P(I)P(II)P(III)=\dfrac{1}{2}(\dfrac{1}{3})(\dfrac{1}{3})=P(I\cap II\cap III)"

 The events are independent.